# Examining Algorithmic Curation on Social Media: An Empirical Audit of Reddit’s r/popular Feed

### v1

@@ 5 RQ2: Rank on r/popular
In this section, we describe our analyses of the factors associated with changes to a post’s rank on r/popular. As seen in Figure 2, a post does not gradually change between nearby ranks over time. Instead, the typical rank trajectory of a post consists of two parts: (1) where the post stays at one rank for several minutes, and (2) where the post suddenly jumps to another, distant rank. To characterize this discontinuous behavior, we used two separate regression models to determine which factors affect: (1) when a post changes rank, and (2) to which rank it changes when it does.
Model 2.1: Multinomial Regression
@@
Model 2.2: Ordinal Regression
The second model is an ordinal regression that estimated the post’s next rank. Only snapshots where a post changed rank were considered so that we could get a more detailed picture of which rank a post changed to when it moved. We interpreted this model as a latent-variable model, where a post’s rank is determined by a continuous latent variable (), and the observed rank depends on which cut points falls between. The model estimated both the cut points and the association between each feature and the expected value of . The flexibility of the cut points within the model makes it well suited to handle the uneven “distances” between ranks—for e.g., it might be less likely for a post to move from rank 2 to rank 1 than it is from rank 99 to 98. We implemented this model in a Bayesian framework with strongly informative priors on the cut points and weakly informative priors on the coefficients. We provided more details on the priors in Appendix A.
@@ Results
Starting with the multinomial regression model, Table 3 shows how a 2 times increase to a feature, excluding the proportion of upvotes, affected the probability of a post moving up or down the feed. Additionally, Figure 3 shows the baseline probabilities for each rank movement across the ranks on the feed. Note that the probability of moving up and down do not sum to 100% because the remaining amount corresponds to the probability that the post will not move in the next snapshot.
@@
Impact of Rank on Movement. We found that posts closer to the top of the feed moved less frequently and less far. Figure 3 shows that the probability that a post moves in either direction reduces as its rank gets closer to 1, and Figure 4 that shows the gaps between cut points close to rank 1 are wider than cut points closer to rank 100. This is consistent with the intuition that there would be less competition toward the top of the feed as it would be rare for posts to achieve that level of “quality.”
@@
Impact of Overall Activity on Rank. Similar to our results in RQ1, the number of recent comments and recent votes were strong factors that increased the probability that a post moved up in rank, as seen in Table 3. In terms of how far a post moved, Table 4 indicates that the total number of comments, recent comments, and recent votes had a similar magnitude effect. However, Table 3 indicates that the total number of comments tended to “stabilize” the rank of a post, decreasing the odds of movement in either direction, but still favoring upward movement. Similarly, the number of subscribers did not seem to increase the probability of upward movement directly, but it still favored upward movement by decreasing the probability of downward movement. These findings are consistent with an engagement-based ranking algorithm, but they highlight the influence of recent activity.
Impact of Undesired Activity on Rank. Table 3 indicates that the proportion of undesired comments slightly decreased the likelihood of moving down the feed, and Table 4 indicates that undesired comments had relatively little impact on where a post jumped in rank to. Thus, while undesired comments had little impact on moving posts up the feed, it reduced the rate at which posts move down.
For reasoning about score’s effect on rank, see Section 8.
@@ 6 RQ3: Engagment on r/popular
Our analyses from RQ1 and RQ2 provided empirical insights into the factors associated with ranking on r/popular. Finally, for RQ3, we investigated how rank on the r/popular feed influences the engagement on posts, specifically the frequency of comments gained before the next snapshot.
Model 3: Negative Binomial Regression
Our goal was to investigate the impact of features in Section 3 on commenting rate, and further, how they impacted the rate of undesired comments differently. We measured commenting rate as the number of comments gained between two consecutive snapshots. For example, if a post had 100 comments during a snapshot and at the next snapshot it had 120, then it has gained 20 comments. Since snapshots are taken approximately 2 minutes apart, this would correspond to a commenting rate of 10 comments per minute.
Since we wanted to investigate if there are differences in how features are associated with undesired commenting rate, we first consider the non-undesired commenting rate to be the reference which we compared the undesired commenting rate against. We used a single negative binomial regression model that estimated both non-undesired and undesired commenting rates using a dummy variable, which indicates that we are estimating undesired commenting rates. Thus, the model directly estimated the association between the features and the non-undesired commenting rate, and differences between how features are associated with undesired compared to non-undesired commenting rate are captured in interaction terms with the dummy variable. In other words, these interaction terms allowed us to inspect whether any of the features from Section 3 or the placement on r/popular has differing effects on undesired comments versus non-undesired comments.
@@ Results
Table 5 shows how our feature set from Section 3, excluding rank which is visualized in Figure 5, is associated with the rate of non-undesired comments, the reference, the rate of and undesired comments.
Impact of Engagement on Non-Undesired Activity. From this table, we found that the number of comments, recent commenting activity, and recent voting activity had the strongest associations with future non-undesired activity. Additionally, the impact of recent commenting activity (61.74%) on non-undesired commenting rate was greater than the one found with the total number of comments (22.33%). This hints that the active discussions played a more important role to future engagement than the total number of discussions that have already occurred.
Impact of Engagement on Undesired Activity. For undesired activity, we found that the proportion of undesired comments had the strongest association (84.43%) out of all the ones found in Table 5. This suggests that having a greater proportion of undesired activity would invite similarly undesired activity in the future.
Impact of Rank on Commenting Activity. Regarding rank’s associations with commenting rate, we found that the rate of non-undesired comments dropped fairly consistently from ranks 2 through 80, visualized in the first subplot in Figure 5. Afterward, the rate of non-undesired comments fell precipitously below rank 80.
Impact of Rank on Undesired Activity. From the second subplot in Figure 5, we observed that the ratio of undesired comments fell as the post is placed lower on the feed. Specifically, the ratio of undesired comments to non-undesired comments fell at a rate of 0.064% per rank (). Additionally, ranks 2 and 3 may indicate that there is a sharp rise in the ratio at the peak of the r/popular feed, however, it is difficult to be sure given the confidence intervals. Overall, we found that posts higher on the r/popular feed had a slightly greater tendency to attract undesired comments compared to posts below it. It is also important to note that the direction of causation is unclear as we previously found that undesired comments may have a slight impact on rank as well.
@@ 7 Discussion
Implications for Content Consumption.
We found that the order in which content is ranked can influence the levels and types of user engagement within algorithmically curated feeds. Similarly, the types of content that are recommended and where they are positioned can affect users’ content consumption practices. Prior work has examined this impact through users’ trust in Google’s ranking algorithm, which leads to a greater number of clicks on top-ranked links regardless of their accuracy or relevance to the query (Joachims et al. 2017). This trust bias that Joachims et al. (2017) refer to, coupled with the increased engagement levels observed in our analysis (see Section 6), can catalyze the spread of misinformation—especially if higher ranks within feeds are misconstrued as an indicator of content quality and trustworthiness. To alleviate some of these pitfalls, some platforms have provided explanations for their algorithmic recommendations, however, Mousavi, Gummadi, and Zannettou (2024) found that those on TikTok in particular were generic and inapplicable to the user. Further research is needed to qualitatively understand how algorithmic rank affects users’ perceptions of content highlighted on social media feeds.
@@
Prior work has shown that sudden popularity can be disruptive to moderation teams that have to manage increased activity, particularly from newcomers (Kiene, Monroy-Hernández, and Hill 2016; Chan et al. 2024). In addition to increased activity levels, we found that posts ranked higher on r/popular corresponded to higher proportions of undesirable activity, which can make handling popularity even more challenging. Despite how integral algorithmic curation is to social media platforms, moderators currently have little influence on algorithmic curation systems. This lack of agency impedes their ability to direct their communities and highlight desirable content—an idea that has been explored in prior work (Choi et al. 2024). Given that moderators have little influence on these systems, there is a need for design interventions that provide moderators with more controls or ways to “override” algorithmic curation systems. These additional affordances can be used to highlight desirable content in feeds, which has theoretical foundations (Kraut and Resnick 2012) and is currently done with existing mechanisms such as awards, upvotes, pins, and flairs (Lambert, Choi, and Chandrasekharan 2024). These additional controls may take the form of various sliders that adjust latent weights on the feed to emphasize different priorities (e.g., increase user diversity on the feed), visual indicators for moderator-assigned contributions, or specific feeds that filter for moderator-selected contributions—similar to the ones the New York Times has to filter for reader- and editor-selected comments (Wang and Diakopoulos 2022).
@@
Despite the pervasiveness of algorithmic curation on social media, studying it as external researchers is exceedingly difficult. This difficulty is compounded by data access restrictions as platforms lock down their APIs, e.g., X/Twitter (Calma 2023) and Reddit (Reddit 2023). The reduction in API accessibility has led researchers to use other methods like sock-puppet accounts (Perriam, Birkbak, and Freeman 2020; Bartley et al. 2021; Bandy and Diakopoulos 2021; Liu, Wu, and Resnick 2024; Mousavi, Gummadi, and Zannettou 2024) and calls for data donations (Jhaver et al. 2023; Chouaki et al. 2024)—each with their own set of drawbacks. These restrictions have led to the reliance on externally-maintained datasets like the now defunct Pushshift—which has been heavily used in prior research (Jhaver, Bruckman, and Gilbert 2019; Atcheson, Koshy, and Karahalios 2024; Kou et al. 2024; August et al. 2024). To avoid these challenges, we developed a robust pipeline to collect large-scale high-fidelity snapshots of Reddit’s trending feed r/popular which can be adapted for other platforms with similar feeds.
However, we recognize that these approaches are extremely time and resource-intensive, posing significant challenges for scalability without adequate financial backing and collaborative support. Although partnerships between academia and industry have the potential to yield important research outcomes, such collaborations are increasingly rare. Given these circumstances, and echoing a theme identified in a 2024 CCC Workshop Visioning Report (Eslami et al. 2024), this presents an opportunity for researchers to share auditing infrastructures and data-sharing protocols that support the collection, storage, and access of social media data for research purposes. For example, in the industry circuit, data clean rooms (Herbrich 2022) are emerging for several companies to have a shared and secure data infrastructure. However, the creation of these infrastructures must also adhere to relevant privacy and legal standards such as GDPR (GDRP 2023) or CCPA (CCPA 2018), especially for data that contain sensitive information that would need to be anonymized. This would enable more research like ours, informing future regulation, while safeguarding the rights and privacy of social media users.
8 Limitations
@@
Our findings are inherently tied to the demographics of Reddit users. Therefore, any attempt to apply our conclusions to other platforms should carefully consider the demographic similarities between Reddit and the platform in question.
The relatively low our models exhibited indicate that there are other factors outside of the ones we have included that influence the outcomes. One such factor may include details about surrounding posts on the r/popular feed. A hint that our model is incomplete is the observation that increasing the score on a post corresponds to a drastic decrease in the odds of staying on the feed, as found in RQ1 and RQ2. This counterintuitive result may be due to a correlation between a post’s score and some unobserved influence. Regardless, future work can build upon these models by including more factors about surrounding posts that are competing for the same finite number of spots on these feeds. Furthermore, future work could also employ more advanced statistical methods like a stochastic transitivity model (Johnson and Kuhn 2013) or a causal framework used in prior work (Chandrasekharan et al. 2017; Saha, Chandrasekharan, and De Choudhury 2019; Jhaver, Rathi, and Saha 2024). These causal approaches will help examine the causal relationship between the different factors that influence and are influenced by algorithmic ranking.
Lastly, because our findings were based on observational data, future work will have to examine more user-centered effects of algorithmic ranking—similar to Salganik, Dodds, and Watts (2006); Joachims et al. (2017), but on social media. Specifically, future work can structure controlled experiments that test the impacts of algorithmic rank on individual users and their perceptions of highly ranked content. These experiments can additionally include interviews or surveys to assess affective characteristics like trust, engagement, and perceived fairness. Together, these studies can offer complementary insights that provide a more comprehensive understanding of algorithmic ranking on social media platforms.

### v3

@@ 6 RQ2: Rank on r/popular
In this section, we describe what factors are associated with a post’s rank on r/popular. As seen in Figure 2, posts do not gradually change between nearby ranks over time. Instead, the typical rank trajectory of a post consists of two parts: (1) where the post stays at one rank for several minutes, and (2) where the post suddenly jumps to another, distant rank. To characterize this discontinuous behavior, we used two separate regression models to determine which factors affect: (1) when a post changes rank, and (2) how far it moves when it does.
Multinomial Regression (Rank Movement)
@@
Ordinal Regression (Rank)
We also aimed to estimate how far a post moves when it does, so we employed a second model: an ordinal regression that estimated the post’s next rank. Only snapshots where a post changed rank were considered so that we could get a more detailed picture of which rank a post changed to when it moved. We constructed a latent-variable model, where a post’s rank is determined by a continuous latent variable (), and the observed rank depends on which cut points falls between. The model estimated both the cut points and the association between each feature and the expected value of . The flexibility of the cut points within the model makes it well suited to handle the uneven “distances” between ranks—e.g., it might be less likely for a post to move from rank 2 to rank 1 than it is from rank 99 to 98. We implemented this model in a Bayesian framework with strongly informative priors on the cut points and weakly informative priors on the coefficients. We provided more details on the priors in Appendix A.
Causal Inference Analysis
We examined the causal relationship between commenting activity and rank movement using an approach similar to the earlier multinomial regression. We used a set of multinomial regression models on our stratified data to estimate the impacts of recent commenting activity (TC) and undesired activity (TU) on the odds of upward movement and downward movement against no movement in the next snapshot.
@@ Results
Starting with the multinomial regression, Table 3 shows the associations between a twofold increase to a feature—excluding the proportion of upvotes—and the probability of a post moving up or down the feed. Additionally, Figure 3 shows the baseline probabilities for each rank movement across ranks. Note that the probability of moving up and down do not sum to 100% as the remainder reflects the likelihood that the post remains in the same position in the next snapshot.
@@
Association Between Rank & Movement. We found that posts closer to the top of the feed moved less frequently and less far. Figure 3 shows that the probability that a post moves in either direction reduces as its rank gets closer to 1, and Figure 4 that shows the gaps between cut points close to rank 1 are wider than cut points closer to rank 100. This is consistent with the intuition that there would be less competition toward the top of the feed as it would be rare for posts to achieve that level of “quality.”
@@
Impact of Overall Activity on Rank. Similar to our results in RQ1, the number of recent comments and recent votes are strong factors associated with increases in the probability moving up in rank, as seen in Table 3. In terms of how far a post moves, Table 4 indicates that the total number of comments, recent comments, and recent votes have a similar magnitude relationship. However, Table 3 shows that the total number of comments tends to “stabilize” the rank of a post, decreasing the odds of movement in either direction, but still favoring upward movement. Similarly, the number of subscribers is not associated with an increase in probability of upward movement directly, but it still favored upward movement by decreasing the probability of downward movement (-2.937%). These findings are consistent with an engagement-based ranking algorithm, but they highlight the influence of recent activity.
Our causal analysis found that increasing the number of comments caused posts to move up the feed by simultaneously increasing the odds of moving up the feed and decreasing the odds of moving down the feed. Specifically, we found strong evidence that the receiving more than 11 recent comments (TC-W) had a significant effect on the odds of moving upward (0.03% in ), and a 99.98% probability () that this effect is positive (mean=10.63%, CI[8.65%, 12.52%]). We also found strong evidence that receiving more than 10 recent comments (TC-W) had a significant effect on the odds of moving downward (0.03% in ), and a 99.98% probability () that this effect is negative (mean=-11.57%, CI[-13.67%, -9.79%]).
We found similar results with the strong treatment (TC-S), though with slightly lower confidence due to the relative lack of suitable strata (see Appendix B). Specifically, we found evidence that receiving at least 24 recent comments (TC-S) had a significant effect (4.53% in ) on the odds of moving upward, and a 99.58% probability () that this effect is positive (mean=11.40%, CI[3.15%, 19.48%]). We also found strong evidence that the same treatment (TC-S) had a significant effect on the odds of moving downward (0.1% in ROPE), and a 99.98% probability () that this effect is negative (mean=-17.80%, CI[-24.19%, -10.60%]).
Impact of Undesired Activity on Rank. Table 3 shows that the proportion of undesired comments slightly associated with a decrease in the likelihood of moving down the feed, and Table 4 indicates that undesired comments has a relatively small but positive relationship with how far a post moves up the feed (negative in terms of ). This is in contrast with Table 3 that only identifies a negative relationship between undesired comments and posts moving downwards.
However, from our causal analysis, we found little evidence that the proportion of undesired comments has any significant effect on rank movement up or down, with a probability in of 97.98% or greater for both the weak (TU-W) and strong versions of the treatment (TU-S).
@@ 7 RQ3: Engagment on r/popular
Our analyses from RQ1 and RQ2 provided empirical insights into the factors are associated with rank on r/popular. Finally, for RQ3, we investigated how rank on the r/popular feed influences the engagement on posts, specifically the frequency of comments gained before the next snapshot.
Negative Binomial Regression
To estimate the relationship between each feature in Section 3 and the comment rate on a post during the next snapshot—measured as the number of new comments between two snapshots—we trained a single negative binomial regression to estimate both the rate of non-undesired comments and undesired comments by using an interaction term that indicates whether the model is estimating the former or the latter. Thus, by default, the model measures the association between each feature and the rate of non-undesired comments and the interaction term measures whether there is any difference in the relationship when estimating the rate of undesired comments instead.
Causal Inference Analysis
To estimate whether moving up the feed causally increases the rate of comments overall, undesired comments, or the proportion of undesired comments, we trained a set of negative binomial regression models. These models utilize our stratified data to estimate the impact moving up the feed one rank (TR-W) and five ranks (TR-S)—i.e., the 75th percentile—on commenting activity. More specifically, we measure each treatment’s effect on the overall commenting rate, undesired commenting rate, and proportion of undesired comments during the period before the next snapshot.
@@ Results
Table 5 shows how our feature set from Section 3—excluding rank which is visualized in Figure 5—is associated with the rate of non-undesired comments and the rate of undesired comments.
Association Between Engagement & Non-Undesired Activity. From Table 5, we found that the number of comments, recent commenting activity, and recent voting activity have the strongest associations with future non-undesired activity. Additionally, the relationship recent commenting activity has with non-undesired commenting rate (61.74%) is greater than the one found with the comments overall (22.33%). This hints that the active discussions plays a more important role to future engagement than the total number of discussions that have already occurred.
Association Between Engagement & Undesired Activity. For undesired activity, we found that the proportion of undesired comments has the strongest association (84.43%) out of all the ones found in Table 5. This potentially suggests that having a greater proportion of undesired activity invites similarly undesired activity in the future.
Impact of Rank on Commenting Activity. Regarding rank’s associations with commenting rate, we found that the rate of non-undesired comments falls fairly consistently from ranks 2 through 80, visualized in the first subplot in Figure 5. Afterward, the rate of non-undesired comments falls precipitously below rank 80.
From our causal analysis, we found strong evidence that moving up the feed increases the comment rate experienced by the post. Both our weak (TR-W) and strong (TR-S) treatments increase commenting rate. Specifically, the weak treatment (i.e., moving up the feed at least one rank) had significant (0.03% in ) mean effect of 8.00% (CI[5.76%, 10.41%]) with a 99.98% probability () that this effect is positive. Similarly, for our strong treatment (i.e., moving up the feed at least five ranks) had a significant (0.05% in ) mean effect of 10.19% (CI[7.14%, 13.09%]) with a 99.98% probability () that this effect is positive. Thus, we found strong evidence that higher positions on the feed incur more comments.
Impact of Rank on Undesired Activity. From the second subplot in Figure 5, we observed that the ratio of undesired comments falls as the post is placed lower on the feed. Specifically, the ratio of undesired comments to non-undesired comments falls at a rate of 0.064% per rank (). Additionally, ranks 2 and 3 may indicate that there is a sharp rise in the ratio at the peak of the r/popular feed, however, it is difficult to be sure given the confidence intervals.
Although Figure 5 shows a slight increase in the proportion of undesired comments at higher ranks, our causal analysis found no disproportionate increase in undesired comments relative to overall commenting. Both treatments had a statistically significant (0.1% in ) effect on the number of undesired comments: 8.65% mean for TR-W (CI[6.29%, 11.18%]) and 10.74% mean for TR-S (CI[7.57%, 13.77%]). However, these effect sizes are comparable to those observed for overall commenting. When analyzing the proportion of undesired comments—thus controlling for the overall comment volume—we found no significant effect (99.98% in ). This suggests that the increase in undesired comments likely coincides with a broader increase in overall engagement, rather than indicating a targeted or disproportionate effect.
@@ 8 Discussion
Implications for Content Consumption. We found that the order in which content is ranked can influence the levels and types of user engagement within algorithmically curated feeds. Similarly, the types of content that are recommended and where they are positioned can affect users’ content consumption practices (Piccardi et al. 2024; Poudel and Weninger 2024). Prior work has examined this impact through users’ trust in Google’s ranking algorithm, which leads to a greater number of clicks on top-ranked links regardless of their accuracy or relevance to the query (Joachims et al. 2017). This trust bias that Joachims et al. (2017) refer to, coupled with the increased engagement levels observed in our analysis (see Section 7), can catalyze the spread of misinformation—especially if higher ranks within feeds are misconstrued as an indicator of content quality and trustworthiness. To alleviate some of these pitfalls, some platforms have provided explanations for their algorithmic recommendations, however, Mousavi, Gummadi, and Zannettou (2024) found that those on TikTok in particular were generic and inapplicable to the user. Further research is needed to empirically understand how algorithmic rank affects users’ perceptions of content highlighted on social media feeds.
@@
Prior work has shown that sudden popularity can be disruptive to moderation teams that have to manage increased activity, particularly from newcomers (Kiene, Monroy-Hernández, and Hill 2016; Chan et al. 2022, 2024). Despite how integral algorithmic curation is to social media platforms, moderators currently have little influence on algorithmic curation systems. This lack of agency impedes their ability to direct their communities and highlight desirable content—an idea that has been explored in prior work (Choi et al. 2025). Given that moderators have little influence on these systems, there is a need for design interventions that provide moderators with more controls or ways to “override” algorithmic curation systems. These additional affordances can be used to highlight desirable content in feeds, which has theoretical foundations (Kraut and Resnick 2012) and is currently done with existing mechanisms such as awards, upvotes, pins, and flairs (Lambert, Choi, and Chandrasekharan 2024). These additional controls may take the form of various sliders that adjust latent weights on the feed to emphasize different priorities (e.g., increase user diversity on the feed), visual indicators for moderator-assigned contributions, or specific feeds that filter for moderator-selected contributions—similar to the ones the New York Times has to filter for reader- and editor-selected comments (Wang and Diakopoulos 2022).
@@
Despite the pervasiveness of algorithmic curation on social media, studying it as external researchers is exceedingly difficult. This difficulty is compounded by data access restrictions as platforms lock down their APIs, e.g., X/Twitter (Calma 2023) and Reddit (Reddit 2023). The reduction in API accessibility has led researchers to use other methods like sock-puppet accounts (Perriam, Birkbak, and Freeman 2020; Bartley et al. 2021; Bandy and Diakopoulos 2021; Liu, Wu, and Resnick 2024; Mousavi, Gummadi, and Zannettou 2024) and data donations (Jhaver et al. 2023; Chouaki et al. 2024) that each have their own set of drawbacks. These restrictions have led to the reliance on externally-maintained datasets like the now defunct Pushshift—which has been heavily used in prior research (Jhaver, Bruckman, and Gilbert 2019; Atcheson, Koshy, and Karahalios 2024; Kou et al. 2024; August et al. 2024). To avoid these challenges, we developed a robust pipeline to collect large-scale high-fidelity snapshots of Reddit’s trending feed r/popular which can be adapted for other platforms with similar feeds.
However, we recognize that these approaches are extremely time- and resource-intensive, posing significant challenges for scalability without adequate financial backing and collaborative support. Although partnerships between academia and industry have yielded important research outcomes, such collaborations are increasingly rare. Given these circumstances, and echoing a theme identified in a 2024 CCC Workshop Visioning Report (Eslami et al. 2024), this presents an opportunity for researchers to share auditing infrastructures and data-sharing protocols that support the collection, storage, and access of social media data for research purposes. For example, in the industry circuit, data clean rooms (Herbrich 2022) are emerging for several companies to have a shared and secure data infrastructure. However, the creation of these infrastructures must also adhere to relevant privacy and legal standards such as GDPR (GDRP 2023) or CCPA (CCPA 2018), especially for sensitive data. This would enable more research like ours, informing future regulation, while safeguarding the rights and privacy of social media users.
Limitations & Future Directions
@@
Our findings are shaped by the demographics of Reddit users. Applying them elsewhere requires careful attention to demographic similarities between Reddit and the platform in question. Our study employed a robust causal inference design that mitigated potential confounds and led to stronger inferences than purely correlational methods. However, we cannot cannot claim “true causality,” due to the lack of “true counterfactuals” and the potential for unobserved confounders. Future research can build on our work through controlled experimental studies to further validate and extend these insights.
The relatively low our models exhibited indicate that there are other factors outside of the ones we have included that influence the outcomes. One such factor may include details about surrounding posts on the r/popular feed. A hint that our model is incomplete is the observation that increasing the score on a post corresponds to a drastic decrease in the odds of staying on the feed, as found in RQ1 and RQ2. This counterintuitive result may be due to a correlation between a post’s score and some unobserved influence. Regardless, future work can build upon these models by including more factors about surrounding posts that are competing for the same finite number of spots on these feeds. Furthermore, future work could also employ more advanced statistical methods like a stochastic transitivity model (Johnson and Kuhn 2013).
Lastly, because our findings were based on observational data, future work will have to examine more user-centered effects of algorithmic ranking—similar to Salganik, Dodds, and Watts (2006); Joachims et al. (2017), but on social media. Specifically, future work could design controlled experiments to examine how algorithmic rank influences individual users and shapes their perceptions of highly ranked content. These studies can also incorporate interviews or surveys to investigate user perceptions of content quality and relevance, as well as affective responses (e.g., trust), and how these factors relate to user engagement and online content consumption. Together, these studies can offer complementary insights that provide a more comprehensive understanding of algorithmic ranking on social media platforms.